The Max Co. has just paid a cash dividend of $3.20 per share. Its current stock price is $75.60 per share. The company is a constant 8% growth firm. What must be the expected rate of return on the company's stock?
Answer #1

Given for The Max Co.

last dividend D0 = $3.20

current stock price, P0 = $75.60

growth, g = 8%

So, next year dividend D1 = D0*(1+g) = 3.2*1.08 = $3.46

Let expected rate of return on the company's stock be Ke.

We know that, using constant dividend growth model, price is calculated using

P0 = D1/(Ke-g)

So, 75.60 = 3.46/(Ke - 0.08)

=> Ke = 12.57%

expected rate of return on the company's stock is 12.57%